A literature professor decides to give a 15 question true- false quiz. If a student takes this quiz just by blind guessing, what the probability that he/she could get exactly 10 answers correctly?
Answer #1

Please note nCx = n! / [(n-x)!*x!]

Binomial Probability = nCx * (p)x * (q)n-x, where n = number of trials and x is the number of successes.

Since there are only 2 options, P(Right) = 0.5 and P(Wrong) = 0.5

Therefore n = 15, p = 0.5, q = 1 – p = 0.5 .

P(X = 10) = 15C10 * (0.5)10 * (0.5)15-10 = 3003 * (0.5)15 = 0.0916